OpenStudy (anonymous):

find all solutions. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

rearrange the equation to equal 0 and let sin(x) = s 7s^2 - 14s +7 =0 you can divide through by 7 s^2-2s+1=0 then factorize s=-1 (repeated) sin(x)=-1 solve for x

OpenStudy (anonymous):

I dont get it, because when i make it = to zero i get 7sin^2x-9sinx+7=0 then i cant divide by 7 through it..

OpenStudy (anonymous):

okay, the -5 only affect the '2' because the '-14sinx' has the sin(x) attached to it

OpenStudy (anonymous):

in the same way you don't alter the sin^2(x) bit

OpenStudy (anonymous):

so would i just subtract 2 then divide by 7? after that i would get sin^2 x-14sinx=-1 then what

OpenStudy (anonymous):

you also need to divide the 14sin(x) so you would have sin^2(x)-2sin(x)=-1 so sin^2(x)-2sin(x)+1=0

OpenStudy (anonymous):

oh okay that makes sense. then whats the next step to find the solutions?

OpenStudy (anonymous):

i just need to know what the next steps are to finish the problem

OpenStudy (anonymous):

@Haseeb96 can you help?

OpenStudy (anonymous):

I told you at the top of this thread! ^^

OpenStudy (anonymous):

You just solve sin(x)=-1 therefore x=sin^-1 (-1) = x

OpenStudy (haseeb96):

since, sin2x=2sinxcosx 7(2sinxcosx)-14sinx+2=-5 14sinxcosx-14sinx+2=-5 14sinx(cosx-1)=-5-2 sinx(cosx-1)=-7\14 sinx(cosx-1)=-1\2 either , sinx=-1\2 or cosx-1=-1\2 sinx=-1\2 or cosx=-1\2 +1 sinx =-1\2 or cosx=1\2 \[x=\sin^{-1} \frac{ -1 }{ 2 } or x=\cos^{-1} \frac{ 1 }{ 2 }\] since, sin 210 = -1\2 and cos 60=1\2 x=240 or x= 60 \[x=\frac{ 4\pi }{ 6 } or x=\frac{ \pi }{ 3 }\]

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

oh I'm sorry, I thought it was sin(^2)x not sin(2x) I apologise! The problem with denoting things differently. My bad